What is the ratio of Kinetic Energy of He and N2 having equal mass and at temperature 300 K and 700 K? 

Grade:12

1 Answers

Harishwar IIT Roorkee
9 years ago
From the kinetic energy formula it can be shown that

\displaystyle \frac {K} {3 N} = \frac {k_B T} {2}